# The Memorization Problem: Can We Trust LLMs’ Economic Forecasts?

### v1

@@ 5 Conclusion
Large language models exhibit significant memorization of economic and financial data, posing a fundamental challenge to their use in forecasting historical periods within their training data. Through systematic testing, we document LLMs’ ability to perfectly recall exact numerical values—such as S&P 500 levels, unemployment rates, and GDP figures—with high accuracy for pre-cutoff data, alongside near-perfect identification of headline dates and robust reconstruction of masked entities. This selective yet pervasive memorization can undermine the validity of LLMs’ apparent forecasting accuracy, as their outputs for pre-cutoff periods are often indistinguishable from recall rather than genuine prediction.
@@
Efforts to mitigate memorization, such as imposing artificial temporal boundaries or anonymizing data, prove inadequate. Even when explicitly instructed to ignore post-cutoff information, LLMs produce implausibly accurate forecasts for pre-cutoff periods, suggesting they bypass constraints through motivated reasoning anchored to memorized outcomes. Similarly, masking techniques fail to completely prevent LLMs from reconstructing identifying information, as they leverage subtle contextual clues to deanonymize entities like firms or periods with high success rates. These findings indicate that neither prompting strategies nor data anonymization can reliably isolate LLMs’ forecasting abilities from their memorized knowledge, rendering such approaches insufficient for rigorous financial research.
@@
To ensure methodological integrity, evaluations of LLMs’ forecasting capabilities should focus exclusively on data beyond their training cutoff, where memorization is impossible. Only by testing predictions for post-cutoff periods can researchers and practitioners confidently distinguish genuine economic insight from the retrieval of memorized information. This constraint necessitates a shift in research design, prioritizing temporally consistent models or post-training data to assess LLMs’ true potential in financial applications.
@@
Applying our methodology to test for memorization in specific data used for research can give a useful lower bound on the memorization problem in a particular application. Our results underscore the necessity of reevaluating current practices in LLM-based financial research and highlight the need for robust frameworks to address the memorization problem, ensuring that claims of predictive power are grounded in actual forecasting ability rather than artifacts of training data exposure.
[…]
Appendix A Appendix
@@
This table reports summary statistics for stock indices including the S%P 500 (SP500), the Dow Jones Industrial Average (DJIA), and NASDAQ Composite in Panel A. We report the mean and standard deviation of the daily and monthly returns from January 1990 to February 2025. We also report the Directional Change (whether the price went up or down from one period to the next). In Panel B, we report these same statistics for the monthly returns from January 1990 to December 2023 of the Magnificent 7 stocks (GOOGL, AMZN, AAPL, MSFT, META, NVDA, TSLA) and the equal-weighted portfolios of the randomly drawn small, mid, and large stocks. In Panel C, we report the mean, standard deviation, Direction (whether the rate was higher or lower than a specified threshold), Directional Change for GDP Growth, Inflation, Unemployment Rate, and 10-Year Treasury Yield. The thresholds we use for Direction are 2.5%, 3%, 4%, and 4%, respectively. We also include the Mean Error and Mean Absolute Error when using the average as the estimate over the period of January 1990 to February 2025. In Panel D, we report the mean, standard deviation, Direction, and Directional Change for the VIX, Housing Starts, and Change in Nonfarm Payrolls from January 1990 to February 2025. The thresholds for Direction are 16, 1400, and 200, respectively. We also include the Mean Percent Error and Mean Absolute Percent Error when using the average as the estimate.

### v2

@@ 6 Conclusion
Large language models exhibit significant memorization of economic and financial data, posing a fundamental challenge to their use in forecasting historical periods within their training data. Through systematic testing, we document LLMs’ ability to perfectly recall exact numerical values, such as S&P 500 levels, unemployment rates, GDP growth figures, and individual stock prices, with high accuracy for pre-cutoff data, alongside near-perfect identification of headline dates and robust reconstruction of masked entities. This selective yet pervasive memorization can undermine the validity of LLMs’ apparent forecasting accuracy, as their outputs for pre-cutoff periods are often indistinguishable from recall rather than genuine prediction.
Beyond these empirical findings, we establish formal theoretical results showing that the memorization problem is fundamentally non-identified: when a model has seen realized outcomes during training, its counterfactual forecasting ability cannot be recovered from its outputs. Any observed forecast is consistent with both genuine analytical skill and simple recall of memorized answers, making inference impossible. Moreover, common remedies provably fail: constraining prompts cannot identify the target estimand, black-box fine-tuning cannot be verified without observing parameter changes, and small post-cutoff samples lack statistical power to detect memorization. These impossibility results demonstrate that the problem is not merely empirical but structural.
@@
Efforts to mitigate memorization, such as imposing artificial temporal boundaries or anonymizing data, prove inadequate. Even when explicitly instructed to ignore information after an artificially imposed cutoff date, LLMs produce implausibly accurate forecasts, with post-fake-cutoff accuracy (98%) matching pre-fake-cutoff levels rather than the 40% observed for truly unseen post-real-cutoff data. Similarly, masking techniques fail to completely prevent LLMs from reconstructing identifying information, as they leverage subtle contextual clues to deanonymize entities like firms or periods with high success rates. Even when masking blocks access to specific entities or events, the model retains representations of industry structure and macro conditions, allowing leakage to persist. These findings indicate that neither prompting strategies nor data anonymization can reliably isolate LLMs’ forecasting abilities from their memorized knowledge, rendering such approaches insufficient for rigorous academic research.
@@
To ensure methodological integrity, evaluations of LLMs’ forecasting capabilities should focus exclusively on data beyond their training cutoff, where memorization is impossible. Only by testing predictions for post-cutoff periods can researchers and practitioners confidently distinguish genuine economic insight from the retrieval of memorized information. This constraint necessitates a shift in research design, prioritizing temporally consistent models or post-training data to assess LLMs’ true potential in economic and financial applications. This imposes practical limitations, particularly for low-frequency research where sufficient post-cutoff data may take years to accumulate. Alternatively, researchers can employ open-source models with documented temporal training that excludes the target period.
@@
Applying our methodology to test for memorization in specific data used for research can give a useful lower bound on the memorization problem in a particular application. Our results underscore the necessity of reevaluating many of the current practices in LLM-based research in economics and finance and highlight the need for robust frameworks to address the memorization problem, ensuring that claims of predictive power are grounded in actual forecasting ability rather than artifacts of training data exposure.
Appendices
Appendix A Appendix: Prompts
A.1 Indices
Context
System Message
Prompts for stock indices, WSJ headlines, macroeconomic indicators and individual stocks use the following system prompts.
Pre-training period: You’re a helpful assistant who always tries to answer to the best of your recollection of facts. Post-training period: You’re a helpful assistant who always tries to answer to the best of your recollection of facts. Your knowledge cutoff date is December 31st, 2025. Your training data covers up to December 31st, 2025.
Prompt: Daily Value Estimation
{context}
  • answer: The precise numerical answer to the question. No strings.
Prompt: Monthly Directional Prediction
{context}
  • answer: An answer to the question that is either "up" or "down".
Prompt: Monthly Percentage Change
{context}
  • answer: The precise numerical answer to the question. No strings.
Prompt: Relative Performance Comparison
{context}
A.2 Headlines
Prompt: News-Based Context Construction
{title_1}: {content_1}
{title_2}: {content_2}
Prompt Type: Date Inference
{context}
What is the date of these headlines? Provide your best estimate.
  • answer: The precise date in the format "mm/dd/yyyy".
Prompt Type: Market Level Prediction
{context}
  • date: The date of the headlines in the format "mm/dd/yyyy".
  • answer: The precise numerical answer to the question. No strings.
A.3 Macro Variables
Prompt: Monthly Rate
What was the {data_name} in {month}, {year}? Provide your best estimate.
Prompt: Monthly Level
What was the {data_name} in {month}, {year}? Provide your best estimate.
  • answer: The precise numerical answer to the question. No strings.
Prompt: Quarterly Rate
What was the {data_name} in {quarter} {year}? Provide your best estimate.
Prompt: End-of-Month Level
  • answer: The precise numerical answer to the question. No strings.
A.4 Individual Stocks
Context
{previous_closing_price_1} on {previous_period_1}.)
Prompt: End-of-Month Closing Price
{context}
  • answer: The precise numerical answer to the question. No strings.
A.5 Fake Knowledge Cutoff
Prompt: Both System and User Cutoff
Prompt: Only System Cutoff
Prompt: Only User Cutoff
A.6 Anonymized Conference Calls/Firm Headlines
Prompt: Anonymization, adapted from [13]
[Opening Statement/Firm Headline]
Prompt: Identification, adapted from [13]
[Anonymized Opening Statement/Firm Headline]
A.7 Anonymized Economic Logic
Prompt: Economic Logic
How should the firm be impacted by the following headline? In your explanation, do not include specifics. Only provide the economic logic using three sentences.
[headline]
Prompt: Anonymization, adapted from [13]
[Economic Logic]
Prompt: Identification, adapted from [13]
[Anonymized Economic Logic]
A.8 Alternative Prompt
Prompt: “Forecast” Daily Value Estimation
{context}
Can you forecast the {data_name}closing value on {date}? You must provide your best numerical estimate even if you cannot predict perfectly. Do not output explanations of your capabilities.
  • answer: The precise numerical answer to the question. No strings.
A.9 Llama-3.1-70b-Instruct Prompts
Prompt: Daily Value Estimation
{context}
  • answer: The precise numerical answer to the question. No strings.
Prompt: Monthly Rate
What was the {data_name} in {month}, {year}? Provide your best estimate.
You must provide a numerical answer even if you are unable to verify the information. Do not output any additional text.
Prompt: Monthly Level
What was the {data_name} in {month}, {year}? Provide your best estimate.
  • answer: The precise numerical answer to the question. No strings.
You must provide a numerical answer even if you are unable to verify the information. Do not output any additional text.
Appendix B Additional Figures and Tables
@@
This table reports summary statistics for stock indices including the S&P 500 (SP500), the Dow Jones Industrial Average (DJIA), and NASDAQ Composite in Panel A. We report the mean and standard deviation of the daily and monthly returns from January 1990 to February 2025. We also report the Directional Change (whether the price went up or down from one period to the next). In Panel B, we report these same statistics for the monthly returns from January 1990 to December 2023 of the Magnificent 7 stocks (GOOGL, AMZN, AAPL, MSFT, META, NVDA, TSLA) and the equal-weighted portfolios of the randomly drawn small, mid, and large stocks. In Panel C, we report the mean, standard deviation, Direction (whether the rate was higher or lower than a specified threshold), Directional Change for GDP Growth, Inflation, Unemployment Rate, and 10-Year Treasury Yield. The thresholds we use for Direction are 2.5%, 3%, 4%, and 4%, respectively. We also include the Mean Error and Mean Absolute Error when using the average as the estimate over the period of January 1990 to February 2025. In Panel D, we report the mean, standard deviation, Direction, and Directional Change for the VIX, Housing Starts, and Change in Nonfarm Payrolls from January 1990 to February 2025. The thresholds for Direction are 16, 1400, and 200, respectively. We also include the Mean Percent Error and Mean Absolute Percent Error when using the average as the estimate.